OpenStudy (anonymous):

write an equation in slope intercept form for the line passing through (1,6) and perpendicular to 3x+7y=1

OpenStudy (compassionate):

When lines are perpendicular that means they have opposite slopes. So, solve the first equation for y 7y = -3x + 1 y = -3/7x 1/7 Now plug in the points (1, 6)

OpenStudy (anonymous):

is it multiplication or addition on the y=-3/7x1/7 i got addition. i may be wrong im just making sure